OpenStudy (ajay1010):

Simplify the expression (sin^3x+cos^3x)/(sinx+cosx)

OpenStudy (anonymous):

the sum of two cubes \(a^3+b^3 \) factors as \[(a+b)(a^2-ab+b^2)\]

OpenStudy (anonymous):

making \[\frac{a^3+b^3}{a+b}=a^2-ab+b^2\] after factoring and cancelling

OpenStudy (ajay1010):

thanks

OpenStudy (anonymous):

yw

OpenStudy (ajay1010):

wait

OpenStudy (anonymous):

To simplify this expression, Am I looking for cos(theta +beta)?

OpenStudy (anonymous):

cos(4x+3x)= cos(4x)cos(3x)- sin(4x)sin(3x) It would be nice. Are we sure of the question?

OpenStudy (anonymous):

and cos(4x + 3x) = cos(4x)cos(3x) - sin(4x)sin(3x) and not the question asked

OpenStudy (anonymous):

The simplest I see is to factor out cos(3x) cos(3x)(cos(4x)-sin(4x)

OpenStudy (whpalmer4):

@Ajay1010 Let \(a = \sin x\) and \(b = \cos x\) then you can rewrite the original problem as \[\frac{\sin^3x + \cos^3x}{\sin x + \cos x} = \frac{a^3+b^3}{a+b} = \frac{\cancel{(a+b)}(a^2-ab+b^2)}{\cancel{a+b}} = a^2-ab+b^2\]Now undo the substitution we made: \[\frac{\sin^3 x + \cos^3 x}{\sin x + \cos x}=(\sin x)^2 - (\sin x)(\cos x) + (\cos x)^2 \]\[=\sin^2 x-\sin x\cos x + \cos^2 x = \sin^2 x + \cos^2 x - \sin x\cos x\] but \(\sin^2 x + \cos^2 x = 1\) so we can further simplify to \[\frac{\sin^3 x + \cos^3 x}{\sin x + \cos x}= 1-\sin x\cos x\] @silveralchemist09 You have misread the problem